If A=15° and c=37, find a.

1
Expert's answer
2022-05-29T17:49:29-0400

Right triangle


"\sin A=\dfrac{a}{c}"

"a=c\sin A=37\sin 15\degree"

"=\dfrac{37(\sqrt{6}-\sqrt{2})}{4}\approx9.5763"